Let f(x,y)= (x + y2 0 you find a 8 >0 such that f(x,y)-f(0,0) <0.01 x,y) = (0,0) (15 marks) whenever
Answer #1

Sol . 12 = ( 344_5227.2 tys-; (219) (2x² + y 2)2 2010) 0 1 6232= (0,0) Now ILF (314) – R10,0)) = 209_522y2+y5.co - t ity?) .= 64 Inl Iso - ab + Jose ² + y 2 / 1 flm,y) – floco)] § 6+ |1n3y2/ so 14/2013) – f100)) coool if 6+ Smitya coool lie. 1 m² 42